Assume that when adults with smartphones are randomly​ selected, 45​% use them in meetings or classes. If 9 adult smartphone users are randomly​ selected, find the probability that exactly 6 of them use their smartphones in meetings or classes.
Answer #1

Solution

Given that ,

p = 0.45

1 - p = 1 - 0.45 = 0.55

n = 9

x = 6

Using binomial probability formula ,

P(X = x) = ((n! / x! (n - x)!) * px * (1 - p)n - x

P(X = 6) = ((9! / 6! (9 - 6)!) * 0.456 * (0.55)9 - 6

= ((9! / 6! (3)!) * 0.456 * (0.55)3

= 0.1160

Probability = 0.1160